# When Does Divide and Conquer Work for Long Context LLM? A Noise Decomposition Framework

### v1

@@ F.2 Predictive Utility for Optimal Chunk Size Estimation
The feasibility of accurately estimating optimal chunk sizes from sparse samples can be understood through our framework’s characterization of model noise. We posit that model noise ($\phi_{model}(x)$) exhibits superlinear growth with context length $L=||x||$. It is reasonable to infer that this underlying noise function $\phi(L)$ is, in practice, often (or can be approximated as) monotonic with respect to $L$ over relevant ranges. Consequently, within the D&C paradigm, reducing the per-worker chunk size is expected to lead to a predictably monotonic (or near-monotonic) decrease in the dominant accumulated model noise component, $\Phi_{model}$.
When $\Phi_{model}$ dominates (and task noise $\Phi_{task}$ is relatively small and aggregator noise $\Phi_{agg}$ is managed), the overall D&C system error as a function of chunk size is likely to exhibit a discernible optimal region (e.g., a convex-like shape or a region where decreasing model noise is balanced by slowly increasing task/aggregator noise) rather than fluctuating randomly and unpredictably. Therefore, a small number of samples across different chunk sizes can provide sufficient information to sketch the rough contour of this error curve, allowing for the identification of the approximate location of this optimal trade-off point without exhaustive evaluation.
Appendix G The impact of overlap in chunking strategies
In our main analysis, we primarily considered non-overlapping chunks. To further explore variations in chunking methodologies, we investigated the impact of introducing a modest overlap between adjacent chunks. This strategy is sometimes employed with the aim of mitigating potential information loss at chunk boundaries, which could otherwise contribute to task noise ($\Phi_{task}$). We conducted experiments using the Llama-70B model on several 128K-token long-context tasks, comparing non-overlapping chunks with chunks having a 1K token overlap (base chunk sizes were 4K and 16K tokens before overlap).
@@ Appendix G The impact of overlap in chunking strategies
The inclusion of a 1K token overlap led to mixed and generally marginal changes in performance. While slight improvements were noted in some specific instances, other cases showed minor degradations or no discernible effect. This suggests that while a small overlap might offer a limited benefit for certain local cross-chunk dependencies, it does not consistently or substantially alter the overall performance trade-offs identified by our framework for these tasks. Moreover, extensive overlap could introduce processing redundancies or even conflicting information for the aggregator, potentially increasing aggregator noise ($\Phi_{agg}$). For the configurations tested, the benefits of this overlap strategy did not appear to consistently outweigh the increased complexity or computational cost.
Appendix H Comparative Analysis with Retrieval-Augmented Generation (RAG)
@@
Appendix I Relation to Architectural Context Extension Methods
@@
Appendix J Connection to effective context length RULER benchmarks
@@
While a longer effective context is generally advantageous, the specific manifestation of model noise ($\Phi_{model}$) and task noise ($\Phi_{task}$) remains highly task-dependent. Complex reasoning or generation tasks can stress long-context capabilities differently than retrieval tasks. These observations reinforce the utility of our noise decomposition framework as an analytical tool. The framework helps explain performance variations by considering the interplay of the task’s intrinsic decomposability, the model’s confusion threshold with increasing context, and the aggregator’s ability to synthesize information, providing insights that extend beyond a single metric of effective context length.
Appendix K Framework implementation in more details

### v2

@@ H.2 Predictive Utility for Optimal Chunk Size Estimation
The feasibility of accurately estimating optimal chunk sizes from sparse samples can be understood through our framework’s characterization of length-induced degradation. We posit that model length-induced degradation exhibits superlinear growth with context length L=xL=||x||. It is reasonable to infer that this underlying length-induced degradation function $g(L)$ is, in practice, often (or can be approximated as) monotonic with respect to $L$ over relevant ranges. Consequently, within the D&C paradigm, reducing the per-worker chunk size is expected to lead to a predictably monotonic (or near-monotonic) decrease in the dominant model term, $\mathcal{L}_{\mathrm{model}}$.
When $\mathcal{L}_{\mathrm{model}}$ dominates (and the task and aggregator terms, $\mathcal{L}_{\mathrm{task}}$ and $\mathcal{L}_{\mathrm{agg}}$, are relatively small or controlled), the overall D&C system error as a function of chunk size is likely to exhibit a discernible optimal region (e.g., a convex-like shape or a region where decreasing length-induced degradation is balanced by slowly increasing decomposition/aggregation overhead) rather than fluctuating randomly and unpredictably. Therefore, a small number of samples across different chunk sizes can provide sufficient information to sketch the rough contour of this error curve, allowing for the identification of the approximate location of this optimal trade-off point without exhaustive evaluation.
Appendix I The impact of overlap in chunking strategies
In our main analysis, we primarily considered non-overlapping chunks. To further explore variations in chunking methodologies, we investigated the impact of introducing a modest overlap between adjacent chunks. This strategy is sometimes employed with the aim of mitigating potential information loss at chunk boundaries, which could otherwise increase cross-chunk dependency effects (the task term, $\mathcal{L}_{\mathrm{task}}$). We conducted experiments using the Llama-70B model on several 128K-token long-context tasks, comparing non-overlapping chunks with chunks having a 1K token overlap (base chunk sizes were 4K and 16K tokens before overlap).
@@ Appendix I The impact of overlap in chunking strategies
The inclusion of a 1K token overlap led to mixed and generally marginal changes in performance. While slight improvements were noted in some specific instances, other cases showed minor degradations or no discernible effect. This suggests that while a small overlap might offer a limited benefit for certain local cross-chunk dependencies, it does not consistently or substantially alter the overall performance trade-offs identified by our framework for these tasks. Moreover, extensive overlap could introduce processing redundancies or even conflicting information for the aggregator, potentially increasing aggregation errors (the aggregator term, $\mathcal{L}_{\mathrm{agg}}$). For the configurations tested, the benefits of this overlap strategy did not appear to consistently outweigh the increased complexity or computational cost.
Appendix J Comparative Analysis with Retrieval-Augmented Generation (RAG)
@@
Appendix K Relation to Architectural Context Extension Methods
@@
Appendix L Connection to effective context length RULER benchmarks
@@
While a longer effective context is generally advantageous, the specific manifestation of the model term ($\mathcal{L}_{\mathrm{model}}$) and the task term ($\mathcal{L}_{\mathrm{task}}$) remains highly task-dependent. Complex reasoning or generation tasks can stress long-context capabilities differently than retrieval tasks. These observations reinforce the utility of our noise decomposition framework as an analytical tool. The framework helps explain performance variations by considering the interplay of the task’s intrinsic decomposability, the model’s confusion threshold with increasing context, and the aggregator’s ability to synthesize information, providing insights that extend beyond a single metric of effective context length.
Appendix M Framework implementation in more details
@@
Appendix N Computational Cost and Latency of D&C vs. Single-Pass
Setup and notation.
Let an input have length $T$ tokens and be split into $n$ equal chunks. Denote
  • $T_{\text{single}}(T)$: latency for a single, large model to process the full input,
  • $T_{\text{dc}}(T/n)$: latency for a smaller worker model to process one chunk,
  • $T_{\text{manager}}(L_{\text{agg}})$: latency for the manager/aggregator to process the short aggregation input of length $L_{\text{agg}}$ (concatenated worker outputs/prompts), where typically $L_{\text{agg}}\ll T$.
We assume worker calls are issued in parallel up to the available concurrency of the deployment.
N.1 Latency
Single-pass wall-clock latency.
$\text{Latency}_{\text{single}}\;=\;T_{\text{single}}(T).$ (7)
D&C wall-clock latency with parallel workers.
$\text{Latency}_{\text{D\&C}}\;\approx\;T_{\text{dc}}(T/n)\;+\;T_{\text{manager}}(L_{\text{agg}}).$ (8)
When is D&C faster?
$\boxed{\,T_{\text{single}}(T)\;>\;T_{\text{dc}}(T/n)\;+\;T_{\text{manager}}(L_{\text{agg}})\,}$ (9)
This inequality is often satisfied in practice because (i) per-call latency grows with input length, so shrinking from $T$ to $T/n$ reduces each worker’s processing time, and (ii) $L_{\text{agg}}$ is short, so the manager adds a small, near-constant overhead. Crucially, D&C does not require $n$ sequential LLM calls; with parallelization the critical path is one worker pass plus a single manager pass.
Practical caveats.
The achievable speedup depends on available parallel capacity (API/cluster concurrency), network overhead, and the manager prompt design. Excessive chunk overlap or verbose worker outputs increase $L_{\text{agg}}$ and can erode gains.
N.2 Monetary cost
Let $p^{\text{big}}_{\text{in/out}}$ be per-token prices for the single, flagship model; $p^{\text{small}}_{\text{in/out}}$ for the smaller worker model; and $p^{\text{mgr}}_{\text{in/out}}$ for the manager. Let $|y|$ be the final output length, $|y_{i}|$ each worker’s output length, and $L_{\text{agg}}$ the manager’s input length.
Single-pass cost.
$\text{Cost}_{\text{single}}\;\approx\;p^{\text{big}}_{\text{in}}\,T\;+\;p^{\text{big}}_{\text{out}}\,|y|.$ (10)
D&C cost.
$\text{Cost}_{\text{D\&C}}\;\approx\;p^{\text{small}}_{\text{in}}\,T\;+\;p^{\text{small}}_{\text{out}}\sum_{i=1}^{n}|y_{i}|\;+\;p^{\text{mgr}}_{\text{in}}\,L_{\text{agg}}\;+\;p^{\text{mgr}}_{\text{out}}\,|y|.$ (11)
Token accounting and implication. With non-overlapping chunks, the dominant input mass remains $T$ in both pipelines. D&C adds only a small extra budget for $L_{\text{agg}}$ and $\sum_{i}|y_{i}|$ (kept short by structured worker outputs). Therefore, when $p^{\text{small}}_{\text{in/out}}\!\ll\!p^{\text{big}}_{\text{in/out}}$ (using compact/open models as workers and a lightweight manager), D&C is typically cheaper while processing roughly the same number of dominant input tokens.
Practical caveats.
Chunk overlap, verbose worker outputs, retries, or using a large manager can increase $L_{\text{agg}}$ and $\sum_{i}|y_{i}|$. These are controlled by (i) minimizing overlap, (ii) constraining worker output schemas, and (iii) keeping the manager focused on short structured inputs.
N.3 Summary of the trade-off
  • Latency: With parallel workers, D&C reduces the critical path from processing $T$ once to processing a single chunk of size $T/n$ plus a short aggregation step (Eq. 9).
  • Cost: Because the dominant token mass $T$ is similar in both methods, shifting those tokens to cheaper worker models and paying only a small aggregation overhead typically lowers total cost.
  • Key condition: Short, structured worker outputs (small $L_{\text{agg}}$) and sufficient parallel capacity make $T_{\text{single}}(T)>T_{\text{dc}}(T/n)+T_{\text{manager}}(L_{\text{agg}})$ and $\text{Cost}_{\text{D\&C}}<\text{Cost}_{\text{single}}$ realistic for long inputs.
These considerations complement the performance-centric analysis in the main paper: when model noise grows with context length and cross-chunk synergy is moderate, D&C can simultaneously improve quality, reduce wall-clock time, and lower cost relative to a single-pass large-model run.
Appendix O Why Static Equal-Length Chunking?
For clear analysis and diagnosis, we adopt a control-variable design: we set the chunk size equal to the per-chunk context length and use equal-length, non-overlapping splits. This way, the performance change as chunk size varies can be primarily attributed to the length effect (model noise), rather than multiple moving parts at once.
Why not use adaptive segmentation in the main experiments?
Adaptive policies (variable lengths, semantic boundaries, routing) simultaneously change:
  • the length distribution across chunks (affecting how model noise scales with length),
  • the boundary placement/routing relative to dependencies (affecting task noise),
  • the structure of worker outputs (affecting aggregator noise).
When several factors vary together, the diagnostic curve (performance vs. chunk size) becomes hard to interpret, and we can no longer cleanly read off the relationship between per-chunk length and model confusion, nor the relative contribution of the three noise sources.
Benefits of equal-length splitting (control of variables)
Equal-length, non-overlapping splits concentrate the degree of freedom onto the per-chunk length $\ell$, which:
  • enables a direct estimation/visualization of how model noise depends on $\ell$ (the basis for our read of Fig. 2/3),
  • stabilizes comparisons between D&C and single-shot across different $\ell$,
  • makes the diagnosis of “model-noise dominated” vs. “task-noise dominated” regimes more reliable.
Empirically, we also observed that small overlaps only yield marginal gains in our setting, suggesting the conclusions are not artifacts of brittle boundaries.
The framework does not preclude adaptivity
Our decomposition naturally accommodates adaptive segmentation; when diagnostics indicate high task noise (strong cross-chunk dependencies), adaptive strategies can be useful in engineering practice. However, the focus of this paper is first to identify the length-driven effect under controlled conditions. In other words, we do not reject adaptivity; we fix chunk lengths to keep variables controlled so that the analysis remains interpretable and reproducible.